OpenStudy (dragontears1128):

Complete the square on the following quadratic x^2+18x+9=0 what value is added to both sides of the equation. ----------------------------------------------- x^2+10x+6=0 what value is added to both sides of the equation.

OpenStudy (mathstudent55):

What are you trying to do?

OpenStudy (dragontears1128):

i fixed the question

OpenStudy (mathstudent55):

Oh, sorry, I see. You need to complete the square.

OpenStudy (mathstudent55):

When start with \(x^2 + bx = 0\) and you need to complete the square, you need to add a number to both sides.

OpenStudy (mathstudent55):

The way you find the number is to take half the value of \(b\) and square it.

OpenStudy (mathstudent55):

Look at your first problem: \(x^2 + 18x + 9 = 0\) In this case, \(b\), the coefficient of \(x\), is 18. Take half of 18 and square it.

OpenStudy (mathstudent55):

Normally, the fist step would be to subtract the constant term to the right side before adding a number to both sides. \(x^2 + 18x + 9 = 0\) Subtract 9 from both sides: \(x^2 + 18x = -9\) Now what do you need to add to the left side to complete the square? Whatever it is, you need to add that number to both sides of the equation.

OpenStudy (dragontears1128):

so your saying that when i subtract 9 from 18x+9 it would = 9

OpenStudy (dragontears1128):

do i add the 9

OpenStudy (dragontears1128):

i have a test on this stuff so i want to make sure I know what I'm doing exactly

OpenStudy (mathstudent55):